A farmer plans to fence a rectangular grazing area along a river with 340 yards. What is the largest area he can enclose?
Answer by nerdybill(7384)

A farmer plans to fence a rectangular grazing area along a river with 340 yards. What is the largest area he can enclose?
.
Let x = width
and y = length
.
Perimeter
2x + y = 340
y = 340-2x
.
Area = xy
Area = x(340-2x)
Area = 340x-2x^2
Area = -2x^2 + 340x
.
Axis of symmetry:
x = -b/2a = -340/(-4) = 85 yards (width)
.
length:
y = 340-2x = 340-2(85) = 340-170 = 170 yards
.
Area:
xy = 85(170) = 14450 square yards